OpenStudy (yacoub1993):

Need Help Please. Verify the identity. (1 + tan^2u)(1 - sin^2u) = 1

OpenStudy (yacoub1993):

@thomaster Help me Please

OpenStudy (yacoub1993):

@Orion1213

OpenStudy (anonymous):

sin^2(A) +cos^2(A) =1 1+tan^2(A) = sec^2(A) Use these identities to proceed

OpenStudy (anonymous):

1-sin^2x+tan^2x-tan^2x sin^2x=1 1-sin^2x+tan^2x-(sec^2x-1)sin^x=1 1-sin^2x+tan^2x-(1/cos^2x-1)sin^2x=1 1-sin^2x+tan^2x-(1/cos^2x)sin^2x+sin^2x=1 1-sin^2x+tan^2x-tan^2x+sin^2x=1 then tanx and sinx will go it is done ....

OpenStudy (rajat97):

you can write 1+tan^2u=sec^2u and you can write 1-sin^2u=cos^2u and then sec^2u=1/cos^2u so (sec^2u)(cos^2u)=1 hope this is helpful:)